How do you convert this problem fraction, decimals, and percent into two forms I tried to do it but can seem to get it right
19
20

9 -1
16 12

Answer by Mathtut(3670)

you need to put lines to show what is being divided by what. I am guessing here but it will show you the method
:
19/20=.95 which when multiplied by 100 gives percentage
.95(100)=95%
:
9/16=.5625(100)=56.25%
:
-1/12=-.0833(100)=-8.33% - there really isnt a negative percentage it simply means you lost 8.33% of something or you were 8.33% under your goal something along those lines